Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

roller

Expert replies
by ketkoag » Thu Apr 30, 2009 11:16 am
A certain roller coaster has 3 cars, and a passenger is equally likely to ride in any 1 of the
3 cars each time that passenger rides the roller coaster. If a certain passenger is to ride
the roller coaster 3 times, what is the probability that the passenger will ride in each of
the 3 cars?

A. 0
B. 1/9
C. 2/9
D. 1/3
E. 1

please explain how to solve this question.
Join the discussion
Source: — Problem Solving |

Re: roller

by kanha81 » Thu Apr 30, 2009 11:29 am
ketkoag wrote:A certain roller coaster has 3 cars, and a passenger is equally likely to ride in any 1 of the
3 cars each time that passenger rides the roller coaster. If a certain passenger is to ride
the roller coaster 3 times, what is the probability that the passenger will ride in each of
the 3 cars?

A. 0
B. 1/9
C. 2/9
D. 1/3
E. 1

please explain how to solve this question.
I think. Not Sure. [spoiler][E][/spoiler]
Person is equally likely to ride in any of 3 cars and we want to find out probab that person rides in 3 diff. cars in 3 times:
t(1) = 3/3
t(2) = 2/2
(3) = 1/1

t(1) * t(2) * t(3) = 1
Want to Beat GMAT.
Always do what you're afraid to do. Whoooop GMAT
Join the discussion

by mike22629 » Thu Apr 30, 2009 11:48 am
IMO C.

First car can be any of the three so Probability = 1

Second Car can be one of the remaining two so Probability = 2/3

Third Car there is only one left not driven so Probability = 1/3

1*(2/3)*(1/3) = 2/9
Join the discussion

by dumb.doofus » Thu Apr 30, 2009 10:32 pm
I am sure the above are valid solutions and correct. I am not trying to make this problem more complex than what it is, but to me the wholistic solution is as follows. Again, this is just my opinion.

Let the three cars be a, b, c
Probability of riding a car = 1/3
Probability of not riding a car = 1-1/3 = 2/3

Probability that the person rides the first car but not the other two = (1/3)(2/3)(2/3) = 4/27
Similarly probability for riding second car and not other two = 4/27
Similarly probability for riding third car and not other two = 4/27

So total probability = 3*4/27 = 12/27

but we see that there are 6 possible arrangements of the car i.e. abc, acb, bac, bca, cab, cba and there are two possibilities each for the three cars to be first, second or third. So we need to divide the final probability by 2

i.e (12/27)*(1/2) = 2/9 and that's the answer..
One love, one blood, one life. You got to do what you should.
https://dreambigdreamhigh.blocked/
https://gmattoughies.blocked/
Join the discussion

by sam117 » Fri May 18, 2012 9:13 am
dumb.doofus wrote:I am sure the above are valid solutions and correct. I am not trying to make this problem more complex than what it is, but to me the wholistic solution is as follows. Again, this is just my opinion.

Let the three cars be a, b, c
Probability of riding a car = 1/3
Probability of not riding a car = 1-1/3 = 2/3

Probability that the person rides the first car but not the other two = (1/3)(2/3)(2/3) = 4/27
Similarly probability for riding second car and not other two = 4/27
Similarly probability for riding third car and not other two = 4/27

So total probability = 3*4/27 = 12/27

but we see that there are 6 possible arrangements of the car i.e. abc, acb, bac, bca, cab, cba and there are two possibilities each for the three cars to be first, second or third. So we need to divide the final probability by 2

i.e (12/27)*(1/2) = 2/9 and that's the answer..
I don't understand why you divide by 2 and not by 6 for instance, can anybody help? Thanks
Join the discussion